OpenStudy (anonymous):

Find the Coordinates for the equations minimum point y=x^2-4x +3

OpenStudy (anonymous):

can you use calculus?

OpenStudy (anonymous):

This should be posted in the math section, but Complete the Square to get the minimum point

OpenStudy (anonymous):

lmfao opps you are in the math section

OpenStudy (anonymous):

ugh i thought i was in physics

OpenStudy (anonymous):

ha! nope but can you help anyways?

OpenStudy (anonymous):

@ihatephysicslol

OpenStudy (anonymous):

\[y = x^2 - 4x + 3 \] which essentially is, \[y = Ax^2 + Bx \] okay so in order to find the minimum you would have to complete the square... so basically all you do is take your A and B value and make a perfect square. How you do that is by isolating your A and B first so.. \[y= (x2−4x) +3\] then taking your B value... \[Cnew =( \frac{ B }{ 2 }) ^2 \] this will give you your new C value for the isolated part of your equation.. so... \[y=(x2−4x + (Cnew))+3\] With the new C value, if we add it into the equation like this \[y=(x2−4x + Cnew - Cnew)+3\] we are basically adding zero... okay so then to make like easier we try to get rid of the negative Cnew by expanding the bracket and add 3 to it so it looks like this... \[y=(x2−4x+Cnew)+3+(Cnew) \] it looks a little bit weird but once you have an actual number for Cnew it will look normal :p okay now with the equation that we've created inside the bracket, you would simply just factor it and it will become a perfect square and you'll have your final equation which should look a little something like this but with numbers \[y = A(x-h)^2 + k \] where (h,k) is basically (x,y) and your "k" or "y" in this case would be your minimum value. Also note that \[(x-h)^2\]because "h" is a negative inside the bracket it will become a positive when you put it into your (h,k) brackets, and vice versa

OpenStudy (anonymous):

oopps \[y=Ax2+Bx + C \] *** forgot the C in the beginning !!

OpenStudy (anonymous):

\[y=Ax^2+Bx+C\] and the two is a square ... these equation things are so frustrating lol

OpenStudy (anonymous):

\[y=(x^2−4x+Cnew)+3+(-Cnew)\] ahh i forgot the negative sign on the Cnew when we expanded it... omg i'm sorry for all these little mistakes